Find the medpoint of the segments with the given endpoints:
a. (4,-12) and (-6,4)
b. (18,28) and (-7,15)
Find the other endpoint of a line segment of a line segmetn with the given midpoint and endpoint:
a. midpoint (6,-1) endpoint (-9,0)

Answer by checkley75(3666)

x=6+4=10 divide by 2 we get 10/2=5 for the x coordinate
y=12+4=16 divided by 2 we get 16/2=8 for the y coordinate .
thus the mid point is (5,8)
---------------------------------------------------------------------
x=18+7=25 divided by 2 we get 25/2=12.5 for the x coordinate
y=28-15=13 divided by 2 we get 13/2=6.5 for the y coordinate.
thus the mid point is (12.5,6.5)
---------------------------------------------------------------------
x=9+6=15 which is half of the line so we add 15 to 6=21 for the x end point
y=1 which is half of the line so we add -1-1=-2 to the -2 & we get 1 for the y coordinate. (21,-2) is the end point.
-------------------------------------------------------------------
THE PROOF IS IN THE GRAPHING OF THESE PROBLEMS. :-)